# Long-Lived Circular Rydberg Qubits of Alkaline-Earth Atoms in Optical Tweezers

### 19 Jan 2024 v1

@@ I Introduction
Arrays of individually controlled and interacting Rydberg atoms based on optical tweezer technology have recently enabled rapid advances in the development of neutral atom quantum simulators and computers [1]. Prominent examples range from large scale simulation of quantum spin models [2, 3], over the implementation of optimization problems [4, 5], to high-fidelty gate operations in quantum circuits [6, 7], and even demonstrations of key steps towards quantum error correction [8, 9, 10, 11, 12].
@@
In this article, we demonstrate the first tweezer-trapped circular states of alkaline-earth atoms, which in contrast to alkali atoms provide a second optically active electron. Combining the richer low-lying electronic structure of divalent atoms with atom arrays already gave rise to powerful new tools [20, 21, 22], for example for optical clock metrology or neutral atom quantum computing [23, 24, 25, 26]. In contrast to low angular momentum Rydberg states (e.g. $\mathrm{S}$- or $\mathrm{D}$-orbitals), circular states of alkaline-earth atoms allow for ionic core excitation in the absence of rapid autoionization, providing a plethora of unique possibilities, such as direct imaging and cooling of Rydberg atoms or coherent optical manipulation via narrow-line core spectroscopy (the latter was recently demonstrated in an atomic beam experiment [27]). Here, we create very high-$n$ ($n=79$) circular Rydberg states of ${}^{88}$r atoms from an array of optical tweezers and demonstrate coherent control of a qubit encoded in circular states separated by two principal quantum numbers which is driven by a two-photon microwave (MW) transition. We demonstrate trapping of the circular Rydberg atom in the optical tweezer exploiting the dominating ${}^{+}$ core polarizability and analyze the effect of the trapping light on the qubit coherence. We observe lifetimes of the circular Rydberg states as long as 2.55 ms, which is about an order of magnitude longer than the free-space black-body decay at room temperature. The long lifetime is achieved by placing the atoms inside a pair of optically transparent capacitor plates, which suppress the black-body field at microwave frequencies.
@@ II Preparing Circular Rydberg States
Our experiments start with an array of ten optical tweezers at a wavelength of $\lambda=539.91\,\rm{nm}$ and a waist of $564(5)\,\rm{nm}$, which are stochastically loaded with single ${}^{88}$r atoms, cooled close to the motional ground state. (see Fig. 1(a) and (b); for details on tweezer loading, in-trap laser cooling, and parity projection in our setup, see Ref. [28]). The atoms are prepared inside a structure consisting of six electrodes. Four of them form a ring structure and allow us to apply electric fields in the $x$-$y$ plane of the tweezer array. The remaining two plate electrodes are placed below and above this ring structure for controlling the electric field along the $z$-direction (direction of axial tweezer confinement). They are fabricated from glass coated with a thin film ($\approx 200\,\mathrm{nm}$ thickness) of indium tin oxide (ITO) [17], which grants excellent optical access for laser cooling, high-NA tweezer generation, and single-atom imaging.
@@
We prepare circular Rydberg states (CRS) by first exciting the atoms from the ${}^{1}\mathrm{S}_{0}$ ground state to the $n=79,{{}^{1}\mathrm{F}_{3}},m=2$ Rydberg level in the presence of a magnetic field $B=0.40(5)\,\mathrm{G}$ pointing along the $z$-direction (see Fig. 1(a)). During this $t_{\mathrm{Ryd}}=20\,\mathrm{\upmu s}$ long off-resonant three-photon excitation (cf. Fig. 1(e)), the tweezer light is blinked off to prevent light shifts. Subsequently, the electron is transferred into the circular Rydberg orbit |$\ket{79\mathrm{C}}$ ($n=79,l=m=78$) via an adiabatic rapid passage protocol. To this end, we first ramp up the electric field along $z$ from zero to $E=478(1)\,\mathrm{mVcm^{-1}}$ within $t_{\mathrm{S}}=5\,\mathrm{\upmu s}$. At this field the initial Rydberg F-state smoothly attaches to the Stark-shifted manifold of high-$l$ levels, effectively forming an equidistant ladder of states with increasing $m$ up to the circular state. This allows for resonant coupling of all levels in that ladder with a single $\sigma^{+}$-polarized radio-frequency (RF) field. In our setup, we generate this field by applying sinusoidal voltages with frequency $f_{\rm{RF}}=70\,\mathrm{MHz}$ and phase shifts of about 90 degrees between pairs of neighboring electrodes on the four ring electrodes. The adiabatic passage is then driven by slowly (within $t_{\mathrm{ARP}}=20\,\mathrm{\upmu s}$) sweeping the electric field along $z$ further up to $597(1)\,\mathrm{mVcm^{-1}}$ through the RF-induced multi-level avoided crossing (see Appendix A for details). Note that prior to this work, circularization via rapid adiabatic transfer has been explored for Rydberg states of much lower principal quantum number ($n\leq 52$) [29].
@@
For state read-out, we exploit state-selective ramped field ionization (SSFI). The ionization field is applied along the $x$-direction and guides the produced ions towards a microchannel plate (MCP) mounted outside the electrode cage for detection. Fig. 1(d) shows histograms of the ionization fields derived from the time-of-flight to the detector before and after the adiabatic rapid passage. For the data after transfer, we apply a resonant two-photon MW pulse which partially transfers |$\ket{79\mathrm{C}}$ to |$\ket{77\mathrm{C}}$, as discussed in more detail in the next paragraph. SSFI provides powerful means to tell apart low-$l$ and high-$l$ states, but does not give enough resolution to distinguish the target circular Rydberg level from not fully stretched ’elliptical states’ with $l<n-1$. In order to probe the fidelity for preparing |$\ket{79\mathrm{C}}$, we implement coherent driving of a microwave qubit encoded in the circular Rydberg levels |$\ket{79\mathrm{C}}$ and |$\ket{77\mathrm{C}}$, which are well separated in the SSFI signal (cf. Fig. 1(d)), from which the population $p_{77}$ ($p_{79}$) in |$\ket{77\mathrm{C}}$ (|$\ket{79\mathrm{C}}$) is extracted. Coupling the qubit states is achieved by an off-resonant two-photon microwave drive at $f_{\rm{MW}}=13.86\,\mathrm{GHz}$ (see Fig. 1(c)). Moreover, after the adiabatic passage and prior to the MW pulse, the electric field is ramped up to approximately $2\,\mathrm{Vcm^{-1}}$, which shifts transitions from |$\ket{79\mathrm{C}}$ to states other than |$\ket{77\mathrm{C}}$ out of resonance. Recording the population transfer $P=p_{77}/(p_{79}+p_{77})$ from |$\ket{79\mathrm{C}}$ to |$\ket{77\mathrm{C}}$ as a function of MW pulse length $t_{\mathrm{MW}}$ reveals coherent Rabi oscillations as depicted in Fig. 1(f). From the maximum population transfer into |$\ket{77\mathrm{C}}$ after a $\pi$-pulse (i.e. after $t_{\mathrm{MW}}=870(5)\,\mathrm{ns}$), we infer about $\epsilon_{\mathrm{CRS}}\approx 70\,\mathrm{\%}$ preparation efficiency of the |$\ket{79\mathrm{C}}$ circular Rydberg level. Damping of the Rabi oscillations is attributed to small shot-to-shot fluctuations of the microwave amplitude at the position of the atoms. It is well modeled by a Gaussian envelope [6], from which we extract an $\nicefrac{{1}}{{e}}$-time of $60(3)\,\mathrm{\upmu s}$. Note that the Rabi frequency also changed by $\approx 10\,\mathrm{\%}$ between measurement sets due to fluctuations of the output power of the microwave generator. This is accounted for by fitting a function with piecewise independent frequencies in the areas separated by the gray-dashed lines.
@@ III Circular State Qubit Coherence
In a next step, we probe the transverse coherence time of the circular state qubit, |$\ket{77\mathrm{C}}\leftrightarrow\ket{79\mathrm{C}}$, via its free induction decay. An exemplary dataset obtained from a standard Ramsey measurement ($\nicefrac{{\pi}}{{2}}$-pulse, wait time $t_{\mathrm{R}}$, $\nicefrac{{\pi}}{{2}}$-pulse) with a $\nicefrac{{\pi}}{{2}}$-pulse time of $t_{\nicefrac{{\pi}}{{2}}}=227(2)\,\mathrm{ns}$ is shown in Fig. 2(a) (blue circles). A sinusoidal fit of the Ramsey signal with a Gaussian envelope $\propto\exp(-t_{\mathrm{R}}^{2}/2{T_{2}^{*}}^{2})$, reflecting a stochastic dephasing process from fluctuations of the qubit resonance frequency, reveals a reversible coherence time $T_{2}^{*}=43(2)\,\mathrm{\upmu s}$. In fact, during this time, longitudinal decay out of the qubit subspace is comparatively small (gray diamonds in Fig. 2(a)), and decoherence is primarily attributed to the first-order magnetic sensitivity of the transition between the circular states, translating into magnetic field noise of roughly $1.5\,\mathrm{mG}$ in our setup (see Appendix B). From the oscillation frequency of the Ramsey fringes, we deduce a MW (red) detuning of $\Delta=66.5(2)\,\mathrm{kHz}$.
@@ IV Tweezer Trapping
The trapping potential seen by the circular Rydberg atom is a sum of two contributions, as depicted in Fig. 3(a). First, the driven motion of the Rydberg electron in the oscillating laser field results in a spatially dependent ponderomotive energy shift [31]. For circular Rydberg orbitals similar or smaller in size than the waist of the optical tweezer, this yields a repulsive potential expelling the Rydberg atom from the trap focus (blue dashed line in Fig. 3(a)). As a consequence, optical trapping of alkali (circular) Rydberg atoms, i.e. atoms with a single valence electron, requires tailored bottle beam potentials with a light intensity miminum at the center [32, 19]. Notably, for very high-$n$ circular Rydberg states, with a radius that exceeds the trap waist, the ponderomotive potential develops a central minimum even for Gaussian beam tweezers, allowing to pin circular orbits via the ponderomotive force alone. For our trap waist, this ’needle-trap’ effect is expected for circular states with $n\gtrsim 100$ (red dashed line in Fig. 3(a)). Alkaline-earth Rydberg atoms feature additional AC polarizability due to the optically active ionic core, resulting in the second contribution to the net trapping potential. For our tweezer wavelength, this results in a strong attractive potential (dotted line in Fig. 3(a)), which overcomes the repulsive ponderomotive force and readily allows for trapping in a standard Gaussian beam tweezer.
@@
Specifically, the net potential depth $U_{0}$ for our |$\ket{79\mathrm{C}}$ is about $\nicefrac{{1}}{{5}}$ of the potential seen by the ${{}^{1}\mathrm{S}_{0}}$ electronic ground state. To ensure qubit trapping in our experiment, we now switch the tweezer light back on immediately after the optical F-state excitation and set the power $P_{\rm{CRS}}$ in each tweezer spot to a value at least 6 times larger ($\approx 350\,\mathrm{\upmu W}$) than the value set for trapping ${{}^{1}\mathrm{S}_{0}}$. Note that we did not observe any effect of the trapping potential on the circular state preparation efficiency in our measurements. In a first set of experiments, we repeat the Ramsey and spin-echo experiments from above, but now with trapped circular Rydberg atoms. Exemplary Ramsey fringes for three different values of $P_{\rm{CRS}}$ are shown in Fig. 4(a)-(c). We observe that the oscillation frequency of the free induction decay decreases for increasing trap depth. This is attributed to the small differential light shift $\delta_{\mathrm{ls}}$ between the two qubit states |$\ket{79\mathrm{C}}$ and |$\ket{77\mathrm{C}}$. Specifically, the tweezer potential is slightly deeper for |$\ket{79\mathrm{C}}$ due to the reduced ponderomotive potential (cf. solid lines in Fig. 3(a)) . This causes a redshift of the qubit resonance in the trap center with respect to the free-space microwave transition ($\delta_{\mathrm{ls}}<0$), and reduces the effective detuning during $t_{\mathrm{R}}$ for increasing $P_{\rm{CRS}}$.
We also identify a reduction in $T{{}_{2}^{\ast}}$ with increasing tweezer depth (Fig. 4(d)). A comparison with results from a semi-classical simulation of the dephasing dynamics suggests that the decrease in $T{{}_{2}^{\ast}}$ is due to (thermal) motion in the trap (see Appendix C). This motion is dominated by the short trap release in combination with the photon recoil during the $20\,\mathrm{\upmu s}$ long Rydberg F-state excitation. Increasing the currently limited Rydberg laser power and thereby shortening the release duration, should allow us to significantly reduce this source of dephasing. We leave this for further work. Notably, also the trap-induced dephasing can be rephased via spin-echo (see Fig. 4(e)), though with a slight reduction of the irreversible transverse coherence time compared to the free-space scenario (dashed line in Fig. 4(e)).
@@
In a second set of experiements, we measure the differential light shifts of our trapped circular Rydberg qubit over a wider range of tweezer depths. To this end, we slowly (within 50 $\upmu$s) ramp the tweezer power to larger values up to $P_{\rm{CRS}}\approx 5.4\,\mathrm{mW}$ after preparing |$\ket{79\mathrm{C}}$, and then perform in-trap microwave spectroscopy on the |$\ket{77\mathrm{C}}\leftrightarrow\ket{79\mathrm{C}}$ qubit transition. Exemplary spectra are depicted in Fig. 3(c), revealing a clear trap-induced light shift on the microwave resonance $\delta_{\mathrm{ls}}<0$, which is extracted from the center value of Gaussian fits to the data. Fig. 3(d) summarizes results for $\delta_{\mathrm{ls}}$ obtained from microwave spectroscopy (red circles) and trap-induced frequency shifts observed in Ramsey spectroscopy (green triangles) as a function of $P_{\rm{CRS}}$. We find good agreement with ab-initio $n$-dependent trap depth calculations (blue-shaded area). Residual discrepancies are attributed to thermal motion in the trap, for which the atom does not ideally probe the trap bottom.

### 04 Mar 2024 v2

@@ I Introduction
Arrays of individually controlled and interacting Rydberg atoms based on optical tweezer technology have recently enabled rapid advances in the development of neutral atom quantum simulators and computers [1]. Prominent examples range from large scale simulation of quantum spin models [2, 3], over the implementation of optimization problems [4, 5], to high-fidelity gate operations in quantum circuits [6, 7], and even demonstrations of key steps towards quantum error correction [8, 9, 10, 11, 12].
@@
In this article, we demonstrate the first tweezer-trapped circular states of alkaline-earth atoms, which in contrast to alkali atoms provide a second optically active electron. Combining the richer low-lying electronic structure of divalent atoms with atom arrays already gave rise to powerful new tools [20, 21, 22], for example for optical clock metrology or neutral atom quantum computing [23, 24, 25, 26]. In contrast to low angular momentum Rydberg states (e.g. $\mathrm{S}$- or $\mathrm{D}$-orbitals), circular states of alkaline-earth atoms allow for ionic core excitation in the absence of rapid autoionization, providing a plethora of unique possibilities. First, the ion core enables conservative trapping in standard Gaussian beam tweezers, which brings scaling advantages in view of power requirements when compared to the bottle traps needed for alkali-atoms. Second, photon scattering at the broad and narrow core transitions can be exploited for direct laser cooling and imaging of the trapped Rydberg atoms, making use of central manipulation techniques developed for trapped ions. Third, combining microwave control of the Rydberg electron with narrow-line optical core spectroscopy involving the ion’s $\mathrm{D}$-level enables local control and read-out of the circular Rydberg qubit via the quadrupole interaction between the two electrons (the latter was recently demonstrated in an atomic beam experiment [27]).
Here, we create very high-$n$ ($n=79$) circular Rydberg states of ${}^{88}$r atoms from an array of optical tweezers and demonstrate coherent control of a qubit encoded in circular states separated by two principal quantum numbers which is driven by a two-photon microwave (MW) transition. We demonstrate trapping of the circular Rydberg atom in the optical tweezer exploiting the dominating ${}^{+}$ core polarizability and analyze the effect of the trapping light on the qubit coherence. We observe lifetimes of the circular Rydberg states as long as 2.55 ms, which is about an order of magnitude longer than the free-space black-body decay at room temperature. The long lifetime is achieved by placing the atoms inside a pair of optically transparent capacitor plates, which suppress the black-body field at microwave frequencies.
@@ II Preparing Circular Rydberg States
Our experiments start with an array of ten optical tweezers at a wavelength of $\lambda=539.91\,\rm{nm}$ and a waist of $564(5)\,\rm{nm}$, which are stochastically loaded with single ${}^{88}$r atoms, cooled close to the motional ground state. (see Fig. 1(a) and (b); for details on tweezer loading, in-trap laser cooling, and parity projection in our setup, see Ref. [28]). The atoms are prepared inside a structure consisting of six electrodes. Four of them form a ring structure and allow us to apply electric fields in the $x$-$y$ plane of the tweezer array. The remaining two plate electrodes are placed below and above this ring structure for controlling the electric field along the $z$-direction (direction of axial tweezer confinement). They are fabricated from glass coated with a thin film ($\approx 700\,\mathrm{nm}$ thickness) of indium tin oxide (ITO) [17], which grants excellent optical access for laser cooling, high-NA tweezer generation, and single-atom imaging.
@@
We prepare circular Rydberg states (CRS) by first exciting the atoms from the ${}^{1}\mathrm{S}_{0}$ ground state to the $n=79,{{}^{1}\mathrm{F}_{3}},m=2$ Rydberg level in the presence of a magnetic field $B=0.40(5)\,\mathrm{G}$ pointing along the $z$-direction (see Fig. 1(a))  [^1] Throughout this work, the excitation rate into the F-state is low enough that interactions do not play a role. Limited by laser power, we typically create not more than one Rydberg atom per realization.. During this $t_{\mathrm{Ryd}}=20\,\mathrm{\upmu s}$ long off-resonant three-photon excitation (cf. Fig. 1(e)), the tweezer light is turned off to prevent light shifts. Subsequently, the electron is transferred into the circular Rydberg orbit |$\ket{79\mathrm{C}}$ ($n=79,l=m=78$) via an adiabatic rapid passage protocol. To this end, we first ramp up the electric field along $z$ from zero to $E=478(1)\,\mathrm{mVcm^{-1}}$ within $t_{\mathrm{S}}=5\,\mathrm{\upmu s}$. At this field the initial Rydberg F-state smoothly attaches to the Stark-shifted manifold of high-$l$ levels, effectively forming an equidistant ladder of states with increasing $m$ up to the circular state. This allows for resonant coupling of all levels in that ladder with a single $\sigma^{+}$-polarized radio-frequency (RF) field. In our setup, we generate this field by applying sinusoidal voltages with frequency $f_{\rm{RF}}=70\,\mathrm{MHz}$ and phase shifts of about 90 degrees between pairs of neighboring electrodes on the four ring electrodes. The adiabatic passage is then driven by slowly (within $t_{\mathrm{ARP}}=20\,\mathrm{\upmu s}$) sweeping the electric field along $z$ further up to $597(1)\,\mathrm{mVcm^{-1}}$ through the RF-induced multi-level avoided crossing (see Appendix A for details). Note that prior to this work, circularization via rapid adiabatic transfer has been explored for Rydberg states of much lower principal quantum number ($n\leq 52$) [30].
@@
We exploit state-selective ramped field ionization (SSFI) for ensemble-averaged state read-out. The ionization field is applied along the $x$-direction via two of the ring electrodes and guides the produced ions towards a microchannel plate (MCP) mounted outside the electrode cage for detection. Fig. 1(d) shows histograms of the ionization fields derived from the time-of-flight to the detector before and after the adiabatic rapid passage. For the data after transfer, we apply a resonant two-photon MW pulse which partially transfers |$\ket{79\mathrm{C}}$ to |$\ket{77\mathrm{C}}$, as discussed in more detail in the next paragraph. SSFI provides powerful means to tell apart low-$l$ and high-$l$ states, but does not give enough resolution to distinguish the target circular Rydberg level from not fully stretched ’elliptical states’ with $l<n-1$. In order to probe the fidelity for preparing |$\ket{79\mathrm{C}}$, we implement coherent driving of a microwave qubit encoded in the circular Rydberg levels |$\ket{79\mathrm{C}}$ and |$\ket{77\mathrm{C}}$, which are well separated in the SSFI signal (cf. Fig. 1(d)), from which the population $p_{77}$ ($p_{79}$) in |$\ket{77\mathrm{C}}$ (|$\ket{79\mathrm{C}}$) is extracted. Coupling the qubit states is achieved by an off-resonant two-photon microwave drive at $f_{\rm{MW}}=13.86\,\mathrm{GHz}$ (see Fig. 1(c)). Moreover, after the adiabatic passage and prior to the MW pulse, the electric field is ramped up to approximately $2\,\mathrm{Vcm^{-1}}$, which shifts transitions from |$\ket{79\mathrm{C}}$ to states other than |$\ket{77\mathrm{C}}$ out of resonance. Recording the population transfer $P=p_{77}/(p_{79}+p_{77})$ from |$\ket{79\mathrm{C}}$ to |$\ket{77\mathrm{C}}$ as a function of MW pulse length $t_{\mathrm{MW}}$ reveals coherent Rabi oscillations as depicted in Fig. 1(f). From the maximum population transfer into |$\ket{77\mathrm{C}}$ after a $\pi$-pulse (i.e. after $t_{\mathrm{MW}}=870(5)\,\mathrm{ns}$), we infer about $\epsilon_{\mathrm{CRS}}\approx 70\,\mathrm{\%}$ preparation efficiency of the |$\ket{79\mathrm{C}}$ circular Rydberg level. Damping of the Rabi oscillations is attributed to small shot-to-shot fluctuations of the microwave amplitude at the position of the atoms. It is well modeled by a Gaussian envelope [6], from which we extract an $\nicefrac{{1}}{{e}}$-time of $60(3)\,\mathrm{\upmu s}$. Note that the Rabi frequency also changed by $\approx 10\,\mathrm{\%}$ between measurement sets due to fluctuations of the output power of the microwave generator. This is accounted for by fitting a function with piecewise independent frequencies in the areas separated by the gray-dashed lines.
@@ III Circular State Qubit Coherence
In a next step, we probe the transverse coherence time of the circular state qubit, |$\ket{77\mathrm{C}}\leftrightarrow\ket{79\mathrm{C}}$, via its free induction decay. An exemplary dataset obtained from a standard Ramsey measurement ($\nicefrac{{\pi}}{{2}}$-pulse, wait time $t_{\mathrm{R}}$, $\nicefrac{{\pi}}{{2}}$-pulse) with a $\nicefrac{{\pi}}{{2}}$-pulse time of $t_{\nicefrac{{\pi}}{{2}}}=227(2)\,\mathrm{ns}$ is shown in Fig. 2(a) (blue circles). A sinusoidal fit of the Ramsey signal with a Gaussian envelope $\propto\exp(-t_{\mathrm{R}}^{2}/2{T_{2}^{*}}^{2})$, reflecting a stochastic dephasing process from fluctuations of the qubit resonance frequency, reveals a reversible coherence time $T_{2}^{*}=43(2)\,\mathrm{\upmu s}$. In fact, during this time, longitudinal decay out of the qubit subspace is comparatively small (gray diamonds in Fig. 2(a)), and decoherence is primarily attributed to the first-order magnetic and second-order electric sensitivity of the transition between the circular states (see Appendix B for details). From the oscillation frequency of the Ramsey fringes, we deduce a MW (red) detuning of $\Delta=66.5(2)\,\mathrm{kHz}$.
@@ IV Tweezer Trapping
The trapping potential seen by the circular Rydberg atom is a sum of two contributions, as depicted in Fig. 3(a). First, the driven motion of the Rydberg electron in the oscillating laser field results in a spatially dependent ponderomotive energy shift [32]. For circular Rydberg orbitals similar or smaller in size than the waist of the optical tweezer, this yields a repulsive potential expelling the Rydberg atom from the trap focus (blue dashed line in Fig. 3(a)). As a consequence, optical trapping of alkali (circular) Rydberg atoms, i.e. atoms with a single valence electron, requires tailored bottle beam potentials with a light intensity minimum at the center [33, 19]. Notably, for very high-$n$ circular Rydberg states, with a radius that exceeds the trap waist, the ponderomotive potential develops a central minimum even for Gaussian beam tweezers, allowing to pin circular orbits via the ponderomotive force alone. For our trap waist, this ’needle-trap’ effect is expected for circular states with $n\gtrsim 100$ (red dashed line in Fig. 3(a)) [34]. Alkaline-earth Rydberg atoms feature additional AC polarizability due to the optically active ionic core, resulting in the second contribution to the net trapping potential. For our tweezer wavelength, this results in a strong attractive potential (dotted line in Fig. 3(a)), which overcomes the repulsive ponderomotive force and readily allows for trapping in a standard Gaussian beam tweezer.
@@
Specifically, the net potential depth $U_{0}$ for our |$\ket{79\mathrm{C}}$ is about $\nicefrac{{1}}{{5}}$ of the potential seen by the ${{}^{1}\mathrm{S}_{0}}$ electronic ground state. To ensure qubit trapping in our experiment, we now switch the tweezer light back on immediately after the optical F-state excitation and set the power $P_{\rm{CRS}}$ in each tweezer spot to a value at least 6 times larger ($\approx 350\,\mathrm{\upmu W}$) than the value set for trapping ${{}^{1}\mathrm{S}_{0}}$  [^2] Clear signatures of trapping are also apparent from the ion arrival times on the MCP. For hold times up to several milliseconds, we observe a shift and broadening for $P_{\rm{CRS}}=0$, which is absent for the trapped case.. Note that we did not observe any effect of the trapping potential on the circular state preparation efficiency in our measurements.
@@
In a first set of experiments, we measure the differential light shift of our trapped circular Rydberg qubit over a range of tweezer depths. To this end, we slowly (within 50 $\upmu$s) ramp the tweezer power to larger values up to $P_{\rm{CRS}}\approx 5.4\,\mathrm{mW}$ after preparing |$\ket{79\mathrm{C}}$, and then perform in-trap microwave spectroscopy on the |$\ket{77\mathrm{C}}\leftrightarrow\ket{79\mathrm{C}}$ qubit transition. Exemplary spectra are depicted in Fig. 3(c), revealing a clear trap-induced light shift on the microwave resonance $\delta_{\mathrm{ls}}<0$. Specifically, the tweezer potential is slightly deeper for |$\ket{79\mathrm{C}}$ due to the reduced ponderomotive potential (cf. solid lines in Fig. 3(a)). This causes a red shift of the qubit resonance in the trap center with respect to the free-space microwave transition ($\delta_{\mathrm{ls}}<0$). In Fig. 3(d) the light shift which is extracted from the center value of Gaussian fits to the data is shown as a function of $P_{\rm{CRS}}$ (red circles). We find good agreement with ab-initio $n$-dependent trap depth calculations (blue-shaded area). Residual discrepancies are attributed to thermal motion in the trap, for which the atom does not ideally probe the trap bottom.
@@
In a second set of experiments, we repeat the Ramsey and spin-echo experiments from above, but now with trapped circular Rydberg atoms. Exemplary Ramsey fringes for three different values of $P_{\rm{CRS}}$ are shown in Fig. 4(a)-(c). We observe that the oscillation frequency of the free induction decay decreases for increasing trap depth. This is attributed to the differential light shift $\delta_{\mathrm{ls}}$ between the two qubit states |$\ket{79\mathrm{C}}$ and |$\ket{77\mathrm{C}}$, causing a red shift and thereby reducing the effective detuning during $t_{\mathrm{R}}$. The extracted light shifts are added to Fig. 3(d) (green circles), where we find good agreement with the expected values as well as with light shifts extracted from the spectroscopy measurements.
We also identify a reduction in $T{{}_{2}^{\ast}}$ with increasing tweezer depth (Fig. 4(d)). A comparison with results from a semi-classical simulation of the dephasing dynamics suggests that the decrease in $T{{}_{2}^{\ast}}$ is due to (thermal) motion in the trap (see Appendix C). This motion is dominated by the short trap release in combination with the photon recoil during the $20\,\mathrm{\upmu s}$ long Rydberg F-state excitation. Increasing the currently limited Rydberg laser power and thereby shortening the release duration, should allow us to significantly reduce this source of dephasing. Perspectively, one may also compensate the differential light shift on the qubit by combining the Gaussian trap with a second Laguerre-Gaussian beam with radial index $p=0$, which counteracts the light intensity gradient around the circular electron orbit. Interestingly, using a pair of Laguerre-Gaussian modes with azimuthal index opposite in sign would allow for locally driving transitions between distant circular states [14].
Notably, also the trap-induced dephasing can be rephased via spin-echo (see Fig. 4(e)), though with a slight reduction of the irreversible transverse coherence time compared to the free-space scenario (dashed line in Fig. 4(e)).